Find the rate at which the reciprocal of a number changes as the number increases.

1
Expert's answer
2021-11-18T06:50:55-0500

Let the number be x and it's reciprocal be y.

So y = "\frac{1}{x}"

Differentiating with respect to time, t

"\frac{dy}{dt} = \frac{d}{dt}(\frac{1}{x})"

"=>\frac{dy}{dt} = \frac{d}{dx}(\frac{1}{x}).\frac{dx}{dt}"

"=>\frac{dy}{dt} = - \frac{1}{x²}\frac{dx}{dt}"

As the number increases, "\frac{dx}{dt} > 0"

Therefore "\frac{dy}{dt}" < 0

So the reciprocal of the number decreases at the rate of the reciprocal of the square of the number multiplied by the rate at which the number increases.